integral{1+2tanx{tanx+secx})^1/2dx=?

 
 
 
 

Grade:Select Grade

1 Answers

rahul
9 years ago
(1+2tanx.tanx+2tanx.secx)^1/2dx=(1+tan^2x+tan^2x+2tanxsecx)^1/2=(sec^2x+tan^2x+2tanxsecx)=[(secx+tanx)^2]^1/2=secx+tanx
therefore,integral(secx+tanx)=ln(secx+tanx)+sec^2x